### v2

# Standardizing the Measurement of Text Diversity: A Tool and Comparative Analysis

Abstract
The diversity across outputs generated by LLMs shapes perception of their quality and utility. High lexical diversity is often desirable, but there is no standard method to measure this property. Templated answer structures and “canned” responses across different documents are readily noticeable, but difficult to visualize across large corpora. This work aims to standardize measurement of text diversity. Specifically, we empirically investigate the convergent validity of existing scores across English texts, and we release diversity, an open-source Python package [^1] https://pypi.org/project/diversity/ for measuring and extracting repetition in text. We also build a platform [^2] https://ai-templates.app based on diversity for users to interactively explore repetition in text. We find that fast compression algorithms capture information similar to what is measured by slow-to-compute $n$-gram overlap homogeneity scores. Further, a combination of measures—compression ratios, self-repetition of long $n$-grams, and Self-BLEU—are sufficient to report, as they have low mutual correlation with each other.
[…]
Standardizing the Measurement of Text Diversity:
A Tool and Comparative Analysis
Chantal Shaib1thanks: Partial work completed while at Adobe Research.    Joe Barrow3footnotemark:     Jiuding Sun1    Alexa F. Siu2 Byron C. Wallace1  Ani Nenkova2 1Northeastern University, 2Adobe Research, 3Pattern Data shaib.c@northeastern.edu
1 Introduction
Figure 1: Part-of-speech pattern visualization with the diversity app. Users upload a corpus, and the UI returns frequent part-of-speech sequences. Users can navigate through different pattern lengths of $n=(2,10)$.
Figure 2: (A) Users start by uploading their own dataset on the right or by selecting one of the existing demo datasets on the left. Once uploaded, users can (B) interactively visualize part-of-speech patterns in the data, (C) interactively search for exact repeated text matches, and (D) calculate lexical diversity metrics.
Evaluation of LLM-generated texts is typically done with respect to accuracy or factuality, e.g., as measured via entailment (Tang et al., 2023), or text quality aspects such as coherence and fluency (e.g., estimated using LLMs as evaluators as in Liu et al. 2023). For tasks where reference summaries are available, the similarity of generated outputs to these is also often measured (e.g., using ROUGE scores; Lin and Och, 2004). A complementary dimension of model performance is diversity, or how much “boilerplate” content is repeated across LLM outputs. For example, prompted to summarize news articles, Llama-2 (Touvron et al., 2023b) tends to generate text that begins with “The article discusses that…", “The article notes that…". Table 1 illustrates example repetitions from various models over a news summarization dataset.
Token Repetition Text
"The article also notes that…" (41/500 )
"The article also mentions that…" (17/500)
"The article notes that the…" (15/500)
"The article discusses the…" (15/500)
"The article also mentions that…" (28/500)
"The article is about the…" (19/500)
"According to the article, a…" (16/500)
Pattern-Matched Text
DT NN VBZ DT JJ NN (84/500)
"The article discusses the recent debate…"
"The book provides a helpful guide…"
"The article discusses the controversial penalty…"
"The article discusses the illicit market…"
NNP NNP DT JJ NN VBD (37/500)
"Christopher Barry, a 53-year-old man, was…"
"Charles Collins, a 28-year-old man, saved…"
"Damian Parks, a 22-year-old student, went…"
"Lynn Fast, a 21-year-old mother, claimed…"
Table 1: Exact text match and repeated part-of-speech patterns across 500 LLM-generated summaries of news articles from CNN/DM. The number of times the pattern occurs is reported parenthetically.
There is a distinct lack of standardization in reporting diversity in machine learning datasets (Zhao et al., 2024). Evaluating diversity is complicated by a lack of tools enabling quantification and exploration of repetitive text in large corpora. This work primarily aims to standardize such reporting, focusing on text-based corpora. We introduce diversity, an open-source Python package for evaluating diversity.1 In addition to this package, we provide a standalone web-based tool that allows users to visualize repetition in their corpus.2 This UI provides an intuitive, efficient tool for text analysis, allowing users to: (i) View repetitive text and part-of-speech $n$-grams of varying lengths; (ii) Quickly compute diversity metrics, and; (iii) Interactively highlight and match repetition in documents. To facilitate future exploration and measurement of lexical diversity, both package [^3] https://github.com/cshaib/diversity and UI code [^4] https://github.com/cshaib/diversity_demo are available under the Apache 2.0 license.
To validate existing diversity metrics, we analyze them over English language outputs from several LLMs. This identifies a few practical, (mostly) independent scores that characterize repetition. We emphasize text length as an important confounder when assessing diversity, and examine diversity in downstream datasets such as instruction tuning. When measuring lexical diversity, many comparisons remain meaningful when accompanied with information about length, but absent this variable no reliable conclusions can be drawn.
In addition to assessing the information overlap of existing metrics, we show that compression ratio—a measure of document compression relative to original size—is a fast, easy to compute score that is sufficient to capture the information in all token/type ratio related alternatives. However, we also find that compression ratios (and all scores considered) are moderately to strongly correlated with text length, complicating interpretation.
Our contributions are as follows. (1) We introduce diversity, a Python package implementing diversity metrics. (2) We host and release source code to a user interface to explore repetition and diversity. (3) We evaluate the convergent validity of existing lexical diversity metrics and highlight compression ratios as efficient measures of diversity. (4) We demonstrate use cases of our package in NLP research via published work.
2 Related Work
Lack of diversity in text may result from repetition of lengthy strings or owe to subtle distributional patterns (Holtzman et al., 2019; Meister et al., 2022, 2023a). We focus on scores that aim to capture overt repetition across outputs, and leave for future work similar analysis of semantic and structural diversity scores (Bär et al., 2012; Shaib et al., 2024).
Conditional generation tasks such as image captioning and dialog summarization have offered observations regarding the diversity of produced texts. Prior work on both of these tasks has documented that models tend to repeat the same text for different contexts. Li et al. (2016) find that four phrases account for about a third of all turns produced by a conversational agent, and Devlin et al. (2015) report that more than half of automatically generated captions are repeated verbatim for different images. Self-repetition (Salkar et al., 2022), i.e., exact repetition of the same $n$-gram ($n\geq 4$) across different summaries is a practical way of measuring repetition for tasks where generated text is lengthier. In such cases exact matches rarely occur but repetition is common, especially relative to the training data (Wang et al., 2023a).
Above we discussed a variety of metrics but it is unclear which of these to use when, and how to efficiently visualize the implications of lower diversity in tasks that comprise only text. Further, prior work has shown that human judgments of diversity are difficult to reliably collect. Humans tend to implicitly conflate quality of text with its diversity, and it can be difficult to separate content and lexical diversity in these assessments (Tevet and Berant, 2021). Motivated by the lack of visualization platforms and the need to provide better methods for human evaluations of repetition in large datasets, we build an interactive tool that allows users to browse highlighted instances of “lower diversity” text, as shown in Figure 1.
2.1 A Smorgasbord of Text Diversity Scores
Scores used to measure diversity across a corpus of texts derive from two core ideas: Computing average similarity between pairs of outputs produced by the same model for different inputs, and computing variants of token/type ratio. The former are adapted from common approaches to reference-based text generation using standard measures of pairwise similarity; the latter track the diversity of vocabulary measured as the ratio of unique words to total words produced, with outputs from a model concatenated into a single text. We first describe each score, and then present insights regarding their mutual redundancy. We also consider run-times, which are lengthy for some metrics and may render them impractical for analysis of a large datasets. All scores are defined for a set of generated texts $D$, each conditioned on its respective input.
Self-BLEU
The quality of text in machine translation, summarization, and image captioning is often reported in terms of overlap with a reference text. This idea can be adapted to measure diversity across different outputs by using one generated text as a “reference” and measuring the similarity of other outputs against this. Self-BLEU measures similarity between all text pairs in $D$ using BLEU (Zhu et al., 2018). BLEU could be replaced with other similarity scores, e.g., ROUGE-L or BERTScore. These variants are called homogenization scores and have recently been used to compare the diversity of texts produced under several conditions (Padmakumar and He, 2023a).
Homogenization Score (ROUGE-L)
All homogenization scores calculate an aggregate similarity across pairs of examples (Equation 1). Here the similarity score of choice is ROUGE-L Lin and Och 2004, which quantifies overlap in terms of longest common sub-sequences between all pairs of text in a corpus instead of the fixed $n$-gram size used in other ROUGE variants:
$\text{hom}(D)=\frac{1}{|D|-1}\sum_{d,d^{\prime}\in D;\>d\neq d^{\prime}}\text{sim}(d,d^{\prime})$ (1)
Homogenization Score (BERTScore)
This homogenization score uses BERTScore to measure similarity between documents in Equation 1. Unlike the other scores, it does not count the repetition of specific tokens, but instead uses BERT embeddings to (ideally) capture “semantic” similarity beyond verbatim $n$-gram matches.
Self-repetition Score
Self-repetition measures the tendency of LMs to repeat long $n$-grams across different outputs (Salkar et al., 2022).
$\text{SRS}(d)=\log\left(\sum\limits_{i=1}^{k}N_{i}+1\right)$ (2)
Where $k$ is total number of 4-grams in a single document $d\in D$, and $N_{i}$ the number of other summaries in which 4-gram $i$ appears. The final score is the sum of $\text{SRS(}d\text{)}$ divided by the number of documents in the corpus $D$.
Moving Average Token-Type Ratio
The token-type ratio for a text is the unique token count divided by the total count of tokens. Moving Average Token Type Ratios (MATTRs) measures the lexical dynamics across a text which is insensitive to text length. This captures the repetition of a given word in segments of text and does not explicitly account for longer repeated sequences (Covington and McFall, 2010).
NN-Gram Diversity Score
NGD extends the idea of token-type ratio to longer $n$-grams (Padmakumar and He, 2023a; Meister et al., 2023b; Li et al., 2023). It is defined as a ratio of unique $n$-gram counts to all $n$-gram counts:
[…]
Where $D\oplus$ denotes the dataset $D$ concatenated into a single string. We use four as the maximum $n$-gram length. This method captures repeated sequences in addition to single token diversity.
Hypergeometric Distribution D
The probability of text under a Hypergeometric Distribution D (HD-D) is an another measure of lexical diversity (McCarthy and Jarvis, 2010). [^5] For both HD-D and MATTR, we use the implementation provided in the lexical-diversity package (https://pypi.org/project/lexical-diversity/). HD-D does not capture repetition of sub-sequences.
2.2 Compression Ratios for Diversity
Compression Ratios
The diversity scores introduced so far are all a function of the number of repeated substrings across outputs. We use gZip to compress the concatenated text of all outputs generated by a model. The compression ratio is then the ratio between the size of the compressed file to that of the original file. High compression ratios imply more redundancy:
$\text{CR}(D)=\frac{\text{size of }D\oplus}{\text{compressed size of }D\oplus}$ (4)
Part-of-Speech Compression Ratio
To capture repeated syntactic patterns, we also compute compression ratios for part-of-speech (POS) tag sequences. We use the NLTK POS tagger  [^6] https://www.nltk.org/api/nltk.tag.html and the Penn Treebank set of 36 tags.
3 Evaluating Repetition with diversity
3.1 Design of the Diversity Package
The diversity package uses NLTK (Bird and Loper, 2004) and SpaCY Honnibal et al. (2020) to tag text with parts of speech and extract $n$-grams. Users can install the package via pip (assuming Python 3.10+). They can then calculate various diversity metrics over a list of texts as follows:
1from diversity import compression_ratio,
2 homogenization_score,
3 ngram_diversity_score,
4 get_pos
5
6text = [
7"I enjoy walking with my cute dog for the rest of the day, ...",
8"I enjoy walking with my cute dog. The only time I felt ...",
9"I enjoy walking with my cute dog, and I really enjoy running..."
10]
11
12# compression ratios
13cr = compression_ratio(text, ’gzip’)
14cr_pos = compression_ratio(get_pos(text)[1], ’gzip’)
15
16# homogenization scores
17hs_rougel = homogenization_score(text, ’rougel’)
18hs_bert = homogenization_score(text, ’bertscore’)
19self_bleu = homogenization_score(text, ’bleu’)
20
21# other
22self_rep = self_repetition_score(text)
23nds = ngram_diversity_score(text, n=4)
Users can also extract repeated text based on part of speech ngrams with the extract_patterns and match_patterns functions.
1n = 5
2top_n = 100
3extract_patterns(text, n, top_n)
First, the user extracts patterns using the extract_patterns function by specifying the n-gram length to search for, and the top_n most repeated n-grams to return. This returns a Python dictionary where the keys are the part-of-speech n-grams and the values are the raw text n-grams matching those patterns. The pattern matches are based on the frequency seen across the entire dataset, i.e., a part-of-speech pattern is only a pattern if it appears in more than 2 texts in the original input. Default values consider the top 100 part-of-speech patterns (sorted by frequency).
Then, using match_patterns, a user can identify all patterns in a single text from the input:
1idx = 2
2match_patterns(text[idx], patterns)
which returns a list of tuples containing the pattern in the first position, and the matched substring in the second position.
Many diversity metrics require pairwise comparisons. With larger datasets, this can become infeasible to compute (see Appendix B). We implement a few methods to increase efficiency: We cache to store already computed pairs, and on the website we chunk incoming data and compute pattern searches over batches.
3.2 Design of the Web UI
The diversity Web UI functionality is built around the Python package. All front-end design uses Vue.js and Tailwind CSS.
The UI offers the same functionality as the package, however allows users to measure and explore repetition without requiring any Python knowledge. Figure 2 shows the main pages of the site: users can begin by (A) either uploading their own text file for analysis or selecting one of the demo datasets provided on the site. Then, the user is prompted to select one of three types of analyses: either (B) to explore part-of-speech patterns, (C) to explore verbatim repeated text, or (D) to measure various diversity metrics of the dataset. Datasets are processed upon upload, and nothing is stored on the backend server aside from the existing demo datasets.
(B) Templates
The templates tab allows users to explore extracted part-of-speech $n$-grams in their selected dataset. The left-most column displays pattern length of $n=[2,10]$. The user can then scroll through all of the templates, select some or all, and see the highlighted text in the middle panel corresponding to the template. The templates are assigned a colour when selected to indicate the corresponding matched text. The right-most column provides a reference for all the part-of-speech tags from SpaCY.  [^7] https://spacy.io/usage/linguistic-features The default pattern length is set to $n=4$. Other lengths will load when selected.
(C) Exact Match
The exact matches tab allows a user to explore exact text matches in their dataset. The top provides two sliders: the left slider allows the user to set a string length to search for ($n=[2,10]$), and the right the minimum number of documents in which the string must appear ($n=[2,10]$). The minimum document occurrence slider defaults to 2. Once selected, the user can scroll through to see the repeated text in bold, and the full document text in which the string appears, as well as the number of documents.
(D) Diversity Metrics
The diversity metrics tabs reports the recommended metrics from our evaluation: Compression Ratio, POS Compression, Self-BLEU, Self-Repetition, and Homogenization with BERTScore. We display these values alongisde a guide to the metrics on the right-hand side.
3.3 Use-Cases
Our implementation of compression ratios over parts-of-speech and tokens (along with BERTScore, Self-BLEU, and self-repetition) have been used in prior works (by other groups) to evaluate diversity in model evaluation and alignment (Lake et al., 2024; Moon et al., 2024; Fernandez et al., 2024), and for reporting diversity over synthetic datasets (Chang et al., 2024; Hastings et al., 2024). Due to its computational efficiency, compression ratios have also been used as optimization parameters in decoding strategies (Lanchantin et al., 2025).
Shaib et al. (2024) use the pattern analysis in diversity to measure and evaluate the prevalence of syntactic patterns in LLMs. Similarly, Wadhwa et al. (2025) extract part-of-speech patterns in distillation tasks for model attribution. Further, insights from our evaluation of diversity metrics have been used to inform how to report diversity with respect to text length and data sizes (Guo et al., 2023; Hastings et al., 2024).
Overall, prior work has shown how NLP researchers can greatly benefit from open access to diversity as well as the evaluation conducted across the implemented metrics. By releasing a user interface, we broaden use-cases for users who would like to explore patterns and diversity in text datasets without needing to use Python.
4 Platform Evaluation: Comparative Analysis of Diversity Metrics
4.1 Data and Models
We compute diversity scores for the outputs of six instruction tuned models on the CNN/DailyMail (Hermann et al., 2015) and XSUM (Narayan et al., 2018) English news summarization datasets: Llama-2 (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), FlanT5-XXL (Longpre et al., 2023), StableLM (Taori et al., 2023; Chiang et al., 2023; Anand et al., 2023), Mistral (Jiang et al., 2023), and StableBeluga (Touvron et al., 2023b; Mukherjee et al., 2023). [^8] All models—except GPT-4—downloaded from HuggingFace (https://huggingface.co/models). We selected these models to cover a range of availability (open and closed), and architectures (encoder-decoder, decoder-only). [^9] We use prompts for summarization provided by each model, where available. See Appendix C.4. The lengths of texts vary considerably by source, for reference and model-produced text alike, so we also note average lengths when reporting diversity.
Figure 3: Correlations between text diversity scores on CNN/DM. Compression ratio correlates strongly with most other diversity metrics.
5 Text Length as a Confounder
To keep compute time and costs manageable, we randomly sample 500 inputs from CNN/DailyMail and XSUM for analysis. Table 2 reports diversity scores for outputs generated by the six LLMs for these inputs. Table 2 (top) reports scores for human-written texts: The article given as input for summarization, the baseline summary comprising the first three sentences of the news article, and the reference summary. These scores serve as a reference point for the diversity scores of the models.
Model
Avg.
Length
CR
CR: POS
NGD
Self-
Rep. (↓)
Hom.
(R-L) (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
MATTR
HD-D
Article 452.25 2.615 5.544 2.637 6.216 0.118 0.696 0.003 0.837 0.896
Article (Lead 3) 75.87 2.369 5.497 3.041 4.276 0.105 0.686 0 0.856 0.892
Reference 51.78 2.277 5.330 3.164 3.842 0.074 0.683 0 0.875 0.919
StableLM 132.71 2.724 5.940 2.673 4.940 0.126 0.689 0.002 0.792 0.867
Mistral 114.88 2.499 5.621 2.926 4.688 0.123 0.697 0.036 0.831 0.880
Llama-2 106.52 2.543 5.684 2.874 4.159* 0.125 0.694 0.001 0.820 0.873
StableBeluga 91.17 2.452 5.644 3.028 4.467 0.121 0.702 0.047 0.846 0.889
FlanT5 63.84 2.453 5.608 2.939 3.608* 0.084 0.667 0 0.833 0.887
GPT-4 55.4 2.361 5.463 3.124 3.909 0.098 0.684 0.001 0.853 0.891
Table 2: Diversity scores for the CNN/Daily Mail dataset. Arrows indicate direction of more diversity. Values indicating less diversity compared to at least one text source that produces longer human texts are bolded; models with scores that are less diverse than those from a model that produces longer summaries are underlined. An asterisk indicates a model more diverse than a shorter human text.
One would expect that human-authored texts would be more diverse than those produced by LLMs (with the caveat that the texts were scraped from the web, and so may contain HTML and page layout artefacts which might be repetitive (Salkar et al., 2022)). The human texts differ by length and the sources of longer texts appear to be less diverse. The association between the length of the produced texts and their diversity is similarly pronounced in the XSUM dataset, as seen Table 3. Text length as a confounder for diversity has been reported in prior work (Salkar et al., 2022), along with potential methods to adjust for this, e.g., sampling blocks of fixed size (Covington and McFall, 2010).
Table 5 reports correlations between the number of words produced by each model and diversity scores. All scores of the token/type ratio family are highly correlated with length, while the pairwise similarity ones are only moderately correlated. Self-BLEU has low correlation with length.
Model
Avg.
Length
CR
CR: POS
NGD
Self-
Rep. (↓)
Hom.
(R-L) (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
MATTR
HD-D
Article 310.20 2.511 5.555 2.756 5.643 0.110 0.695 0.002 0.838 0.892
Article (Lead-3) 55.94 2.316 5.454 3.107 3.999 0.103 0.683 0 0.860 0.891
Reference 21.04 2.276 5.409 3.211 2.914 0.081 0.673 0 0.877 0.888
StableLM 109.20 2.745 6.008 2.636 4.687 0.130 0.695 0.002 0.78 0.854
Llama-2 102.48 2.634 5.802 2.795 4.618 0.128 0.687 0.002 0.795 0.858
Mistral 95.18 2.531 5.708 2.911 4.495 0.132 0.698 0.044 0.819 0.867
StableBeluga 88.46 2.461 5.673 2.992 4.418 0.124 0.698 0.046 0.837 0.88
GPT-4 62.15 2.394 5.531* 3.079 4.041 0.104 0.682 0 0.848 0.886
FlanT5 20.93 2.666 6.222 2.743 2.868 0.114 0.665 0.001 0.756 0.842
Table 3: Diversity scores for XSUM summaries. Arrow indicate the direction of more diverse texts for each score.
Model
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Article 2.162 5.095 2.719 0.666 0
Article (Lead 3) 2.179 5.093 2.719 0.663 0
Reference 2.230 5.314 2.663 0.667 0
Llama-2 2.345 5.636 2.919 0.663 0.002
GPT-4 2.213 5.425 2.666 0.663 0
FlanT5 2.490 5.737 2.707 0.665 0.001
StableLM 2.342 5.521 2.823 0.664 0.001
Mistral 2.308 5.689 2.736 0.659 0
StableBeluga 2.210 5.436 2.663 0.659 0
Table 4: Diversity metrics for XSUM summaries, with outputs from each model truncated to the length of the shortest. All scores are directly comparable.
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.867 0.832 0.81 0.904 0.875 0.579 0.235 0.79 0.855
Table 5: Correlation between scores and total word counts (concatenated text) for CNN/Daily Mail.
6 Diversity of Model Summaries
The confound of length complicates reporting. On both CNN/DM and XSUM (cf. Tables 2 and 3), StableLM produces the longest summaries. All scores indicate that these are the least diverse, most likely due to the length confound. In both sets of results, we look for models that produce shorter summaries that are less diverse. These findings are notable and hold, despite length differences.
Three types of differences are marked in the tables. Model summaries that are shorter but less diverse than human summaries are marked in bold. Human texts here are written by journalists, so the expectation is that they would be more diverse. More bold entries in a column indicate that the score captures differences between human and machine diversity, a desirable trait. Underlined entries denote models that are less diverse than other models that produce longer summaries. The more underlined entries there are for a model, the more indicators there are that its output is less diverse. Asterisks mark models that appear more diverse than a human text of shorter length.
The most interesting diversity scores are those that capture differences between human and automatically produced text, without necessarily committing to an interpretation of which source is preferable. On the CNN/DM dataset, Hom. (BERT) and MATTR are the two scores that detect no differences between human and model texts. Compression ratio for part of speech sequences is the score that identifies the most differences between human and model-generated text. Self-repetition stands out as the only score that identifies model generated text as more diverse on the CNN/DM dataset. From this analysis, CR:POS and self-repetition emerge as prime candidates of reportable scores, while Hom. BERT is less useful.
7 Correlation Analysis
We present three sets of correlation analyses between (i) different diversity scores, (ii) the same diversity score across datasets, and (iii) diversity scores and standard reference-based evaluations. Despite the large number of diversity scores in our list, they all revolve around $n$-gram repetition. Do these capture different (complementary) information? To assess this we compute the correlations between all pairs of scores, reported in Figure 3.
Compression ratio is highly to moderately correlated with other $n$-gram scores. The only weak correlations are with Self-BLEU and Hom. (BERT). Given the degenerate behavior of Hom. (BERT) on the analysis of summaries, reporting Self-BLEU only is advisable. Finally, self-repetition is only moderately correlated with with other scores, so it is informative to report this as a standard diversity score. Correlations are similar on XSUM summaries (Appendix 6), reinforcing the recommendations here. Diversity analysis on the CNN/DM and XSUM datasets did not indicate consistent system behavior. We report the analysis in Table 10.
Our guiding assumption is that output diversity and self-repetition are aspects of model behavior that are not captured by existing evaluation approaches. Here we directly test this assumption. We compute the system level correlation between the diversity scores and the traditional BERTScore and ROUGE evals, shown in Figure 4. Reference-based evaluations are only weakly correlated with the diversity metrics. Self-BLEU, however, is moderately anti-correlated with with both ROUGE-1 and BERTScore.
Figure 4: Correlations between diversity metrics, BERTScore, and ROUGE-1. Both reference-based metrics are weakly correlated with CR and Hom. (BERT), and moderately anti-correlated with Self-BLEU.
8 Truncating to Control Length
Model
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Article 2.268 5.25 2.763 0.676 0
Article (Lead 3) 2.274 5.25 2.762 0.658 0
Reference 2.189 5.179 2.763 0.674 0
Llama-2 2.96 5.627 2.847 0.674 0.001
GPT-4 2.287 5.376 2.761 0.672 0
FlanT5 2.288 5.389 2.779 0.673 0
StableLM 2.393 5.537 2.884 0.672 0.001
Mistral 2.32 5.415 2.812 0.67 0
StableBeluga 2.288 5.46 2.766 0.671 0
Table 6: Scores on CNN/DM summaries truncated to the shortest summary length for a given input.
For each input for summarization, we truncate all summaries to the length of the shortest one produced by any of the sources as a crude means to remove the influence of length on scores. The resulting scores are directly comparable across sources, listed in Tables 6 and 4. Compression ratio and Self-BLEU scores indicate that model-produced text is less diverse than human text. Hom. (BERT) scores barely vary across sources, further supporting the recommendation that this is not a useful score to report. On the CNN/DM dataset, Self-BLEU indicates that Llama-2 and StableLM are the most repetitive models. Compression ratio also ranks these two models as the least diverse. The results are consistent on XSUM, but for that dataset Flan-T5 is also highly ranked and the most repetitive.
The truncation approach to control for length is not practical for published research or leaderboards. Introducing a new source of texts would require recomputing the scores for other sources one may want to compare with, which is impractical and sometimes impossible when the outputs from other sources are not available. Future research will have to search for more practical alternatives.
9 Discussion and Recommendations
The diversity package and platform provides a useful way to analyze and visualize diversity in datasets. For reporting metrics, our in-depth analyses reveal that compression ratio is an excellent score to report, easy to compute and strongly correlated with other scores used in past work. Compression ratio of part of speech sequences capture differences between human and model-generated text. Self-repetition zeros in only on repetition of longer $n$-grams across generations, and is only moderately correlated with compression ratios. Finally Self-BLEU is only weakly correlated with the previous three, so is a good complement score to report. In our analyses, we identified several drawbacks of BERTScore: it does not show differences between human and model-generated text and barely varies when adjusted for length.
Length of the analyzed text has to be reported alongside all these scores. When length differs, scores are not meaningfully comparable. Truncating and downsampling text is one way to produce a set of results that are intuitively comparable. Different random draws of the sample chosen to represent a dataset will likely differ in diversity; the selection may lead to unwarranted conclusions. Truncating texts prevents any possibility of discovering repetitive behavior towards the end of longer text. Future research into a principled solution for this problem is urgently needed. Despite all this, we were able to glean meaningful insights about differences in diversity between human and model-produced text for summaries, essays and instructions.
Finally, diversity offers a platform and package in which researchers from a variety of domains can use to facilitate evaluation.
10 Limitations
In the work presented here we do not explore human approaches to evaluating the diversity of collections of text. These are straightforward when the produced text is fairly short, as in judging the diversity of a set of questions generated for a given document (Sultan et al., 2020) or the diversity of possible continuation of a conversation (Tevet and Berant, 2021). Longer texts, as in the case of summaries, and larger collections, as in the case of instruction datasets are harder to judge for diversity.
Acknowledgements
We gratefully acknowledge the National Science Foundation (RI 2211954) for supporting this work.
Appendix A Appendix
Appendix B Run-Time Considerations
When analyzing the diversity of large volumes of text, run-time considerations become relevant. Figure 5 provides insights about the feasibility of obtaining scores for large samples [^10] Run on a single NVIDIA Quadro RTX 8000 GPU.. The compression ratio scores are fast, with text compression utilities specifically optimized for speed. Self-repetition takes longer but acceptable time. Self-BLEU and Hom. (BERT) are prohibitively slow.
Figure 5: Mean run time (log-scale) on CNN/DM summaries. Run times increase with the number of text for the analysis. Even for small datasets, Self-BLEU and BERTScore homogenization are slow.
Appendix C Broader Applications
The guiding motivation for this work has been to develop standardized and informed approach to the analysis the diversity of text produced by LLMs. The standardization of scores will facilitate analysis in broader settings. Here we provide two examples: human writing, with and without facilitation from a LLM, and instruction tuning datasets.
Human Story Writing
Padmakumar et al. (2023) presented an analysis of human-written stories, where people wrote either by themselves or with the help of GPT-3 or GPT-3.5 Turbo. They find that using LLMs as writing partners leads to greater homogenization of the stories. As reported by Padmakumar et al. (2023), we find that all diversity scores agree that people writing independently produce the more diverse texts (cf. Table 7). Here, story length is not an issue because the average length of stories in each setting are comparable: 375 words for writing without help, 372 words when writing with GPT-3 and 370 when writing with GPT-3.5.
Dataset
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Solo 2.901 5.314 5.873 0.604 0.018
GPT-3 2.940 5.371 5.911 0.613 0.020
InstructGPT 3.064 5.462 5.966 0.631 0.022
Table 7: Diversity scores over essays. Working with an LLM correlates with lower diversity.
Instruction-tuning Datasets
The quality and diversity of instructions are likely to result in more robust and capable systems (Sanh et al., 2022; Mishra et al., 2022). We analyze the diversity of five instruction-tuning datasets: Open Assistant (Köpf et al., 2024), Super-NaturalInstructions (Wang et al., 2022), Unnatural Instructions (Honovich et al., 2023), Alpaca (Wang et al., 2023b), and Dolly (Conover et al., 2023).  [^11] We provide further details in Appendix C.6.
In Table 8 we report diversity scores. Here datasets are ordered by size; we therefore expect that scores will be sorted in diminishing order in each column. Only deviations from this ordering are reportable. We provide details about the number of instructions and words in Appendix C.6. Open Assistant instructions are remarkably diverse compared to the other datasets, and all diversity scores for it are more favorable than that for other datasets. Unnatural instructions are remarkable in the opposite direction, with outlier scores that are so much higher, they are likely not due to length entirely. We provide an analysis of the diversity scores with the length controlled in Appendix C.7.
Given the large dataset sizes, ranging from 15-80k data points, we do not compute the homogenization scores nor Self-BLEU, as the computation time is infeasible. For approximately 50k instructions, the estimated computation times ranged from 48 to 800 hours for these scores. This case study highlights the relevancy of the run-time analysis for computing score that we presented in the previous section.
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.886 6.731 3.969
Unnatural Instructions 4.191 7.278 9.868
Alpaca 3.119 6.61 3.105
Super-NaturalInstructions 2.675 5.749 3.456
Dolly 2.578 6.214 2.935
Table 8: Diversity scores for instruction datasets. We do not include Self-BLEU nor Hom. (BERT) due to long run times. Datasets are ordered by size and differ vastly in length, so only scores for which a smaller dataset is less diverse are meaningfully interpretable.
C.1 Human Assessments of Diversity
C.2 Examples of Repetitive Patterns
Table 9 show more examples of repeated sentence structures (using part-of-speech tags) from Padmakumar et al. (2023).
Dataset Token Repetition Text Pattern-Matched Text
GPT-3 "In my opinion…" 41/100 PRP VBZ RB JJ TO VB 15/100
"It is also vital to discern…"
"It is very easy to construct…"
"It is largely inappropriate to try…"
"It is morally acceptable to focus…"
PRP VBP IN DT NN IN 12/100
"I don’t like the damsel in…"
"I fear that a cycle of…"
"I feel that an acknowledgement of…"
"I find that the inflection of…"
Instruct-GPT "In my opinion…" 25/100 MD VB DT JJ NN IN 20/100
"It is important to…" 20/100 "…can have a huge variety of…"
"Up with the news…" 15/100 "…can have a negative effect on…"
"…can have a positive impact on…"
"…can have a sturdy framework for…"
PRP VBZ RB JJ TO VB 12/100
"It is also important to realize…"
"It is fairly common to hear…"
"It is indeed surprising to hear…"
"It is probably impossible to keep…"
"…it becomes very cringy to watch…"
Solo "In my opinion…" 22/100 PRP VBZ JJ TO VB IN 9/100
"In my opinion, the…" 13/100 "It is crucial to recognize that…"
"When it comes to…" 11/100 "It is crucial to remember that…"
"In my opinion, I…" 10/100 "It is unjustifiable to assume that…"
"It is important to acknowledge that…"
PRP VBP IN DT JJ NN 10/100
"I believe for the right person…"
"I do on a regular basis."
"I fall into the second group."
"I live in a small town…"
Table 9: Examples of exact text-match and repeating part-of-speech patterns in essays from Padmakumar and He (2023b). The number of times the pattern occurs is shown in parenthesis.
C.3 Correlation Between Metrics
Self-BLEU scores are almost perfectly correlated between the two datasets; they appear to not be affected by text source. The other scores are still moderately to highly correlated but as already observed, models are ranked differently. When reporting diversity, source of analyzed data also has to be taken into account, in addition to length.
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.83 0.695 0.885 0.87 0.841 0.921 0.991 0.799 0.654
Table 10: Score correlations for each text diversity score between the CNN/DM and XSUM datasets.
C.4 Summarization Prompts
Table 11 details the prompts and format used to generate the summaries for the news datasets. We follow the formats recommended provided by each model, and insert the along with the instruction for summarization.
Model Model Size Prompt
Llama-2 7B [TEXT] [INST] Summarize the above text. [/INST]
GPT-4 - [TEXT]. Summarize the above text.
Flan-T5 11B Summarize this article: [TEXT]
StableLM 7B [TEXT] <|<|USER|>|>Summarize the above text.
<|<|ASSISTANT|>|>
Mistral 7B ### Instruction: Summarize the following: ### Input: [TEXT].
### Response:
Table 11: Prompts used for each model to generate a summary. [TEXT] is replaced with the input article.
C.5 XSUM Metrics
Figure 6 shows the correlations between all pairs of metrics for the XSUM dataset. The correlations show that compression ratio is highly to moderately correlated with other n-gram scores, similar to the findings for the CNN/DM dataset.
Figure 6: Correlation table between scores on XSUM.
C.6 Instruction Datasets, Details
Open Assistant
is a collection of crowdsourced instructions (Köpf et al., 2024). The data was collected under detailed guidelines and includes questions that reflect real-life situations.
Super-NaturalInstructions
A corpus comprising crowdsourced instructions that transform 200 benchmarks and intermediate evaluation results into a set of instructions and demonstrations (Wang et al., 2022).
Unnatural Instructions
An (almost) automatically created dataset, using instructions from the SuperNatural-Instructions dataset to automatically generate new instructions (Honovich et al., 2023). To increase diversity, each instruction was also paraphrased. Honovich et al. (2023) compare the diversity of instructions in Unnatural and Super-Natural Instructions with pairwise BERTScore similarities (within each dataset), and find that the similarities are much higher in Super-NaturalInstructions.
Alpaca
This dataset is created following the Self-instruct dataset (Wang et al., 2023b). GPT-3 was prompted to create instructions and demonstrations based on a seed of 175 human-written instructions. Crucially, the collection method includes a diversity filter, only including model-written instructions if their ROUGE-L similarity is less than 0.7 with an existing instruction. Length of instructions and demonstrations is also controlled for as a criterion for inclusion in the final instruction dataset.
Dolly
A set of human instructions and demonstrations, collected by Datrabricks employees (Conover et al., 2023). By design, they cover only eight classes of popular tasks: creative writing, closed and open QA, summarization, information extraction, classification and brainstorming.
Table 12 shows the number of instructions, the typical length of an instruction and average number of words per instruction set. All vary, making it even harder to control for length. Truncating makes less sense here, and down-sampling the number per instructions is counter-productive given our goal to understand the diversity of the entire dataset. We do make use of these instruments given the lack of alternatives, but note that more meaningful solutions are urgently needed.
Dataset # Instructions Avg. # Words Total # Words
Open Assistant 84,437 78.10 6,594,646
Unnatural Instructions 66,010 38.05 2,511,737
Alpaca 52,002 10.06 523,329
Super-NaturalInstructions 4550 92.58 421,228
Dolly 15,011 12.37 185,816
Table 12: Average number of words, and size of the instruction datasets. Numbers correspond to the training set available from Huggingface. For Super-NaturalInstructions, we filter for English-only instructions using the langdetect library.
C.7 Instruction Datasets, Length Controlled
Table 13 shows scores for instructions downsampled to the size of the smallest dataset, and truncated to the length of the shortest instructions in the remaining data. Again, the Open Assistant dataset stand out as most diverse, while the Unnatural Instructions dataset is markedly less diverse than the others. Self-repetition in the related Super-Natural and Unnatural instructions is notably high. The human instructions in Dolly compare favorably with automatic instructions, especially when bearing in mind that only eight tasks are covered in it. CR:POS points to Super-natural instructions as the most diverse. We do not have a convincing explanation of why it compares so favorably against others on this score.
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.370 5.402 1.741
Unnatural Instructions 6.036 8.421 5.595
Alpaca 3.301 6.044 2.020
Super-NaturalInstructions 2.458 1.844 4.859
Dolly 2.832 5.504 2.235
Table 13: Truncated diversity scores for instruction datasets.